# Page Curve and Entanglement Dynamics in an Interacting Fermionic Chain

### v2

@@ I Introduction
Inspired by the black hole information paradox, we model a many-body system in this work where we have an arbitrary bi-partition which we refer to as the “system” and the “environment”. We show that starting from a non-entangled initial state, we obtain a Page curve for the evolution of the vNEE. Analogous to black holes, the system is interacting while the environment is free. The situation where the system is also free is exactly solvable in the thermodynamic limit [27]. The free case has also been analyzed in Refs. [28, 29]. We also study Rényi entropies which all exhibit Page curve behavior. In particular the infinite order Rényi entropy (the so-called min-entropy) develops a non-analyticity in its time evolution suggesting a breakdown of the semi-classical picture. Indeed, we find a quantum phase transition in the corresponding entanglement Hamiltonian where the critical time precedes the Page time. This allows us to draw a quantum phase diagram where the bending down of vNEE belongs to the quantum critical regime above the quantum critical point (see Fig. 7). We emphasize that the construction is analogous to black hole physics but does not contribute directly towards the resolution of the black hole information paradox, albeit some parallels do exist between our construction and the one of black hole information paradox.
The paper is organized as follows. We present the model in Section II where we also present the methodology and the results obtained in this work. We proceed to discuss the entanglement dynamics of the non-equilibrium case in Section III where we explicitly study the entanglement entropy, min-entropy as well as the local particle number density for different parameter values. We find a non-analytic behavior in the min-entropy, which we interpret (using the methodologies developed in Section II and Appendix A) as a quantum phase transition in the corresponding entanglement Hamiltonian. We provide a unified perspective on the time-dependent entanglement Hamiltonian for a model with conserved charge to explain the observed non-analyticity in Section IV. Finally, we extrapolate to the thermodynamic limit via performing a regression analysis in Section V where we observe the persistence of non-analyticity at non-zero time for weak interactions in the system but vanishing for intermediate interactions. We also provide an analytical explanation and justification in Section V for the observed behavior in the strongly interacting limit. We finally conclude in Section VI. The appendices provide further details of the calculation and additional numerical results.
@@ II.1 Results
    The slope of min-entropy is (essentially) linear until the non-analytic point, and this allows us to do a scaling analysis and extrapolate to the thermodynamic limit using linear regression. We find in the thermodynamic limit that for weak interaction ($V/t_{h}<1.0$), there exists a finite, non-zero critical time at which the non-analyticity appears while for intermediate interactions ($V/t_{h}\sim 1.0$), the critical time tends to zero. We also find an acute sensitivity on the coupling strength $g$ between the system and the environment (Tables 1 and 2) where a weaker coupling $g$ requires a stronger interaction for the thermodynamic critical value to vanish.
@@ III Entanglement Dynamics and Page Curve
As discussed above, we present the entanglement dynamics of Rényi entropies in this section for a total system size $L=50$ for model (1). Motivated by the black hole setup, in the following, we have $N\gg M$ while $L=M+N=50$ is fixed. We consider different system sizes, namely $M=3,4,5,6$ and $7$ in order to perform a scaling analysis $M\to\infty$ to the thermodynamic limit (TL), see Section V. Since in all cases treated we consider time scales before any particles reach the right edge of the lattice at site $L$, in this way we indeed obtain the behavior in the TL of the entire set up. The parameter values are fixed in the model Eq. (1) as $t_{s}=t_{e}=1\equiv t_{h}$ and the system and the environment are coupled weakly, in particular $g\in\{0.25,0.50\}$. Two interaction regimes are considered: weak ($V/t_{h}<1.0$) and intermediate ($V/t_{h}\sim 1.0$) whereas an analytical explanation is provided in Section V for the strongly interacting case ($V/t_{h}\gg 1$).
@@
Non-analyticity in min-entropy.— We study the min-entropy that is defined in Eq. (4) and we find a non-analyticity in the temporal evolution at some critical time $t_{c}<$ ${t_{\text{Page }}}$. Fig. 5(a) displays data for $M=5$ for different values of $V$; the behavior for further values of $M$ is presented in Appendix B. As can be seen, several non-analytical points appear as a function of time. In the following, we mainly focus on the first non-analyticity, since it will help us in better understanding the behavior before reaching the maximum of the entanglement. As can be seen in Fig. 5(a), the value of $V$ has a minimal influence on the value of the decayed fraction $1-m(t)/M$, at which the first non-analyticity appears. This illustrates the persistence of the non-analytic behavior also in the presence of interactions. The question arises, if the behavior is universal for all values of the interactions, or if eventually the interactions will lead to different behavior. To study this, we investigate the behavior in the TL. In Fig. 5(b) we now keep $V$ fixed and study the dependence of the critical value of $1-m(t)/M$ on $M$. We see that the kink position happens at smaller values of $1-m(t)/M$ when increasing $M$. This is further studied in Section V, where we extrapolate this value to the TL via linear regression. We refrain from studying the later non-analyticities, since we expect that, similar to the findings in the free case [27], in the TL they will smear out, so that only the first non-analyticity will persist.
@@ IV Time-dependent entanglement Hamiltonian and the conserved charge picture
V Analysis in the thermodynamic limit
@@
We discuss a scaling analysis by performing a linear extrapolation to the thermodynamic limit. Considering that the short time behavior of the min-entropy until the first non-analyticity is effectively linear, we expect this to capture the leading order of the behavior. As motivated by the free case [27], we expect later non-analyticities to smear out in the thermodynamic limit. For a fixed interaction strength $V=0.8$ (as shown in Fig. 5), we select the $x$-coordinates of the first non-analytic points for different system sizes that we denote as $1-m(t_{c})/M$ and perform a linear regression as shown in Fig. 8. The system parameters treated and the results are summarized in Table 1 and a CSV file containing a summary of regression analyses is provided in the research data management [37]. We find an acute sensitivity on the coupling strength $g$ between the system and the environment, where we find that the non-vanishing critical value persists for higher interaction strengths when the coupling is weaker. For both cases of the coupling ($g=0.5$ and $g=0.25$), the value corresponding to the free case $V=0.0$ indeed matches the value obtained in the exactly solvable free model as carried out in Ref. [27]. At finite $V>0$, for the smaller value of $g$ we find a non-vanishing $t_{c}$-value in the TL for $V\lesssim 1$. However, when further increasing $V$, or in the case of $g=0.5$, we always obtain $t_{c}\approx 0$; a conservative estimate of the numerical errors after the extrapolation gives an errorbar of $\sim 3$ times the standard deviation shown, so that we conclude that these values are zero within our obtained numerical accuracy. This is also true at $V=0$, in agreement with Ref. [27] (see Table 1). Note that we have performed regression analyses for the critical value of decay fraction instead of critical time (as plotted in Fig. 3 for the vNEE); the scaling analyses for the critical time is provided in the Appendix B where we observe the same pattern.
@@
We can make this interpretation concrete to show the vanishing of the critical values in the thermodynamic limit. Consider the strongly interacting case where $V/t_{h}\gg 1$ (where $t_{h}\equiv t_{s}=t_{e}$ which has been set to $1.0$ in this work). Then one particle decays to the environment and accordingly a hole enters the system. In this strongly interacting case, no higher order processes will play any relevant role and the system will freeze in this state and no further decay of particles will be observed. We have verified this numerically for $V=3.0$ and $V=5.0$ for time scales accessible to us (see Appendix C for details). Then we can write an ansatz for the number of particles in the system where only two states play a role: (1) when no particle has decayed, and (2) when one particle has decayed into the environment. Since we are interested in finding the crossing of the ground state of the entanglement Hamiltonian and the dynamics is dominated by these two possibilities, only the largest and the second-largest Schmidt coefficients, namely $1-\lambda(t)$ and $\lambda(t)$ respectively, are playing a role such that $\lambda(t=0)=0$ (see Appendix C where we have also provided numerical evidences). The ansatz reads as
@@ VI Summary and Outlook
Along the same lines as in the free case [27], we expect the later non-analyticites (after the first one) to smear out in the TL. Using the linear growth of the min-entropy up to the first non-analyticity, we used linear regression to extrapolate to the thermodynamic limit where we found that there exists a non-zero critical time for weak interactions at which the quantum phase transition in the entanglement Hamiltonian develops. However, as the interaction strength increases, the critical value vanishes in the thermodynamic limit, implying an instantaneous dynamics leading up to the non-analytic behavior (see Fig 10). Note that also for $t_{c}=0$ the semiclassical description loses its applicability, so that, as long as $t_{c}\lesssim{t_{\text{Page }}}$, the proposed scenario keeps its validity (see Fig. 10). Furthermore, we found an acute sensitivity on the coupling strength $g$ between the system and the environment where the weaker the coupling $g$, the stronger the interaction is required for the thermodynamic critical value to vanish. Our analyses show that one must be careful before naively applying perturbation theory by switching on the interactions because of the underlying quantum critical point that might influence the behavior of the system even for finite Rényi entropies. Moreover, the order parameter for the transition, whose schematic is sketched in Fig. 7, is the change in particle number density where a depletion transition drives the quantum phase transition. Near the critical point, the particle density develops a power law behavior with an exponent $1/2$ for the free case [27]. Since the weak interactions are nearly the same as the free case, including the thermodynamic critical value (see Table 1), we strongly expect the power law behavior to stay the same. Further analyses about the details of the phase transition such as the scaling laws remain a work for the future.
@@
Returning to our original motivation from the black hole information paradox, we have drawn analogies between the black hole setup and our lattice setup. Despite the obvious differences, there does exist some similarities between the two. A recent proposal in black hole physics has gathered attention, namely the island formula [39, 40, 41], where a quantum extremal surface develops during the temporal evolution of vNEE right before the Page time. The presence of islands is crucial after the Page time for the reproducibility of the Page curve. However, the emergence of islands is not smooth, thereby leading to a non-analytic behavior in the evolution of vNEE. This suggests the possibility of whether Page curves in the black hole information paradox can be understood via the mechanism highlighted in this work. The question remains about the validity of our claims to black holes, and to which extent we can further understand the emergence of these islands as being related to some form of quantum phase transition in the associated entanglement (equivalently, modular) Hamiltonian.

### v3

@@ I Introduction
Inspired by the black hole information paradox, we model a many-body system in this work where we have an arbitrary bi-partition which we refer to as the “system” and the “environment”. We show that starting from a non-entangled initial state, we obtain a Page curve for the evolution of the vNEE. Analogous to black holes, the system is interacting while the environment is free. The situation where the system is also free is exactly solvable in the thermodynamic limit [27]. The free case has also been analyzed in Refs. [28, 29]. We also study Rényi entropies which all exhibit Page curve behavior where a recent study for local Hamiltonians for one dimensional chains found at most a linear in time behavior [30]. In particular the infinite order Rényi entropy (the so-called min-entropy) develops a non-analyticity in its time evolution suggesting a breakdown of the semi-classical picture. Indeed, we find a quantum phase transition in the corresponding entanglement Hamiltonian where the critical time precedes the Page time. This allows us to draw a quantum phase diagram where the bending down of vNEE belongs to the quantum critical regime above the quantum critical point (see Fig. 7). We emphasize that the construction is analogous to black hole physics but does not contribute directly towards the resolution of the black hole information paradox, albeit some parallels do exist between our construction and the one of black hole information paradox.
The paper is organized as follows. We present the model in Section II where we also present the methodology and the results obtained in this work. We proceed to discuss the entanglement dynamics of the non-equilibrium case in Section III where we explicitly study the entanglement entropy, min-entropy as well as the local particle number density for different parameter values. We find a non-analytic behavior in the min-entropy, which we interpret (using the methodologies developed in Section II and Appendix A) as a quantum phase transition in the corresponding entanglement Hamiltonian. We provide a unified perspective on the time-dependent entanglement Hamiltonian for a model with conserved charge to explain the observed non-analyticity in Section IV. In Section V, we perform finite-size scaling to extract universal finite-size scaling exponents across all interaction strengths. Near the non-analytic point (critical point), we determine the order parameter exponent $\beta$. Its value converges for all interaction strengths to that of the non-interacting limit in the thermodynamic limit, illustrating that the transition is a continuous phase transition and that the system belongs to the same universality class as the free counterpart. Finally, we extrapolate to the thermodynamic limit via performing a regression analysis in Section VI where we observe the persistence of non-analyticity at non-zero time for weak interactions in the system but vanishing for intermediate interactions. We also provide an analytical explanation and justification in Section VI for the observed behavior in the strongly interacting limit. We finally conclude in Section VII. The appendices provide further details of the calculation and additional numerical results.
@@ II.1 Results
    We investigate the nature of the quantum phase transition through finite-size scaling. Figure 8 demonstrates excellent data collapse across all interaction strengths (additional plots are provided in Appendix D), confirming universal finite-size scaling exponents (Table 1). Furthermore, we find a universal order parameter exponent $\beta\sim 0.5$ (Table 2) across all interaction strengths, confirming a continuous quantum phase transition. This value agrees with the thermodynamic limit result $\beta=0.5$ for the non-interacting chain in Ref. [27] (Eq. (61) therein), suggesting the transition belongs to the same universality class.
@@
    The slope of min-entropy is (essentially) linear until the non-analytic point, and this allows us to do a scaling analysis and extrapolate to the thermodynamic limit using linear regression. We find in the thermodynamic limit that for weak interaction ($V/t_{h}<1.0$), there exists a finite, non-zero critical time at which the non-analyticity appears while for intermediate interactions ($V/t_{h}\sim 1.0$), the critical time tends to zero. We also find an acute sensitivity on the coupling strength $g$ between the system and the environment (Tables 3 and 4) where a weaker coupling $g$ requires a stronger interaction for the thermodynamic critical value to vanish.
  5. (v)
@@ III Entanglement Dynamics and Page Curve
As discussed above, we present the entanglement dynamics of Rényi entropies in this section for a total system size $L=50$ for model (1). Motivated by the black hole setup, in the following, we have $N\gg M$ while $L=M+N=50$ is fixed. We consider different system sizes, namely $M=3,4,5,6$ and $7$ in order to perform a scaling analysis $M\to\infty$ to the thermodynamic limit (TL), see Section VI. Since in all cases treated we consider time scales before any particles reach the right edge of the lattice at site $L$, in this way we indeed obtain the behavior in the TL of the entire set up. The parameter values are fixed in the model Eq. (1) as $t_{s}=t_{e}=1\equiv t_{h}$ and the system and the environment are coupled weakly, in particular $g\in\{0.25,0.50\}$. Two interaction regimes are considered: weak ($V/t_{h}<1.0$) and intermediate ($V/t_{h}\sim 1.0$) whereas an analytical explanation is provided in Section VI for the strongly interacting case ($V/t_{h}\gg 1$).
@@
Non-analyticity in min-entropy.— We study the min-entropy that is defined in Eq. (4) and we find a non-analyticity in the temporal evolution at some critical time $t_{c}<$ ${t_{\text{Page }}}$. Fig. 5(a) displays data for $M=5$ for different values of $V$; the behavior for further values of $M$ is presented in Appendix B. As can be seen, several non-analytical points appear as a function of time. In the following, we mainly focus on the first non-analyticity, since it will help us in better understanding the behavior before reaching the maximum of the entanglement. As can be seen in Fig. 5(a), the value of $V$ has a minimal influence on the value of the decayed fraction $1-m(t)/M$, at which the first non-analyticity appears. This illustrates the persistence of the non-analytic behavior also in the presence of interactions. The question arises, if the behavior is universal for all values of the interactions, or if eventually the interactions will lead to different behavior. To study this, we investigate the behavior in the TL. In Fig. 5(b) we now keep $V$ fixed and study the dependence of the critical value of $1-m(t)/M$ on $M$. We see that the kink position happens at smaller values of $1-m(t)/M$ when increasing $M$. This is further studied in Section VI, where we extrapolate this value to the TL via linear regression. We refrain from studying the later non-analyticities, since we expect that, similar to the findings in the free case [27], in the TL they will smear out, so that only the first non-analyticity will persist.
@@ IV Time-dependent entanglement Hamiltonian and the conserved charge picture
V Finite Size Scaling and Critical Exponent
We start by doing a finite size scaling and then evaluating the critical exponent $\beta$ corresponding to the order parameter for the transition, namely the difference of particles in the environment. We first proceed by explaining the scaling relations and then present the results for our case. The parameters are the same as considered throughout this work: $t_{s}=t_{e}=1.0$ as well as the system-environment coupling $g=0.5$. Starting with the critical time $t_{c}$ (which vanishes in the thermoydnamic limit for the system-environment coupling $g=0.5$, as shown in Section VI), we have the following scaling ansatz to illustrate that indeed there is a linear dependence on $1/M$ (thereby justifying the usage of linear regression in Section VI):
$t_{c}\sim M^{-1/c}$ (9)
where $c$ is the scaling ansatz exponent and $M$ is the number of sites in the system. Then once we have obtained the exponent $c$, we have the finite size scaling for the order parameter $(1-m(t)/M)$ which is the decayed fraction of particles to the environment
$(1-m(t)/M)M^{b/c}\sim(t-t_{c})M^{1/c}$ (10)
where we are in the small neighborhood of $t_{c}$. Notice that the transition happens when the first particle leaks into the environment. That’s why the order parameter is zero for $t<t_{c}$ while it becomes non-zero for $t>t_{c}$ in the thermodynamic limit (see Section IV above for a detailed discussion). Using this scaling ansatz, we can extract the scaling ansatz exponent $b$. Finally, the ground-state energy of the entanglement Hamiltonian develops non-analyticity in time and we perform a scaling analysis in the neighborhood of the non-analytic point as
$\Big{(}\frac{S_{\text{min}}}{M}\Big{)}M^{a}\sim(t-t_{c})M^{1/c}$ (11)
which allows us to calculate the scaling ansatz exponent $a$. Once we have performed the scaling, we expect a perfect collapse of the order parameter as well as the min-entropy for various system sizes $M=\{3,4,5,6,7\}$. Indeed, we find universal exponents $a,b$ and $c$ such that for all values of $V=\{0.0,0.4,0.8,1.2\}$, we find a perfect collapse. The free case $V=0.0$ is used to benchmark against the analytical calculations performed in the thermodynamic limit in Ref. [27]. We provide one such example of collapse here for the case of $V=1.2$ in Fig. 8 while other plots for different values of $V$ are provided in Appendix D. We find the values of all three scaling ansatz exponents ($a,b,c$) to be around $1.0$. The precise values are shown in Table 1. This shows a consistent collapse for the order parameter as well as the non-analyticity in the ground state energy of the entanglement Hamiltonian.
Then we find the critical exponent $\beta$ of the order parameter in the small neighborhood of $t_{c}$ such that $t>t_{c}$ which is given by
$\Big{(}1-\frac{m(t)}{M}\Big{)}=A(t-t_{c})^{\beta}\qquad(t>t_{c}).$ (12)
For the free case, in the thermodynamic limit, the critical exponent is found to be $\beta=0.5$ as shown in Eq. (61) of Ref. [27]. For finite systems, we fix the value of $M$ and then perform the scaling to extract the values of $A$ and $\beta$. We know that the prefactor $A$ is not universal, so we restrict our analysis for $\beta$. Then for a given interaction strength $V$ and fixed total number of sites $L$, we aggregate the value of $\beta$ for $M=\{3,4,5,6,7\}$. Within reasonable accuracy, we find the universal critical exponent $\beta$ as the free case [27] for all values of $V$. We provide one such example for $M=5$ and $V=0.8$ in Fig. 9. Aggregated values for various $V$ are summarized in Table 2. Therefore we conclude the robustness of such temporal quantum phase transitions that the interactions don’t change the critical exponent associated with the order parameter .
Figure 8: Top row: raw data in the vicinity of the critical point where top left is for the decayed fraction of particles against time, while the top right plots min-entropy against time (axes are re-scaled system size as shown in the plots). Bottom row: finite-size scaling for the raw plots in the neighborhood of the critical (non-analytic) point, showing a perfect collapse. Using the scaling relations explained in Section V, we find universal values of scaling ansatz exponents $a,b$ and $c$ for all values of $V$. The plot is shown for $V=1.2$, $g=0.5$, $t_{s}=t_{e}=1.0$, $L=50$ and $M=\{3,4,5,6,7\}$. Plots for other values of $V$ are shown in Appendix D. Therefore, we have a consistent collapse for both the order parameter as well as the ground state energy of the entanglement Hamiltonian.
Table 1: Scaling exponents, as explained in Section V, for various values of $V$. Other parameter values are $g=0.5$, $t_{s}=t_{e}=1.0$, $L=50$ and $M=\{3,4,5,6,7\}$.
V $c$ $b$ $a$
$0.0$ $1.01$ $0.99$ $0.99$
$0.4$ $1.02$ $0.99$ $0.98$
$0.8$ $1.02$ $0.98$ $0.97$
$1.2$ $1.02$ $0.99$ $0.98$
Figure 9: We perform the scaling relation in Eq. (12) for $M=4$ and $V=0.8$ to get extract the critical exponent $\beta$ associated to the order parameter. The post-critical time data is used that is in a small neighborhood of the critical time. The post-critical time data is shown for clarity purposes where we expect the scaling relations to fail as they are valid only in the vicinity of the critical time. The black dashed line denotes the actual fit while the blue dashed line denotes the curve in Eq. (12) for $\beta=0.50$ while using the same value of $A$ as obtained via fitting in the black dashed curve. Accordingly, we find to a reasonable accuracy the value of the critical exponent $\beta$ same as that of the free case in the thermodynamic limit as shown in Eq. (61) in Ref. [27]. Other parameters are: $L=50$, $g=0.5$ and $t_{s}=t_{e}=1.0$. We perform an average over different $M=\{3,4,5,6,7\}$ for a fixed $V$ to get aggregated value for finite systems and the results are summarized in Table 2.
Table 2: Critical exponent, as explained in Eq. (12), for various values of $V$. Other parameter values are $g=0.5$, $t_{s}=t_{e}=1.0$ and $L=50$. We have averaged for $M=\{3,4,5,6,7\}$ to get a value for finite systems and within reasonable accuracy, we find $\beta\approx 1/2$ — the same as the free case $V=0.0$ in the thermodynamic limit [27]. Accordingly, interactions don’t affect the robustness of such temporal quantum phase transitions.
V $\beta$ $A$
$0.0$ $0.51$ $0.42$
$0.4$ $0.52$ $0.40$
$0.8$ $0.54$ $0.37$
$1.2$ $0.52$ $0.33$
VI Analysis in the thermodynamic limit
@@
We discuss a scaling analysis by performing a linear extrapolation to the thermodynamic limit (as justified above in Section V). Considering that the short time behavior of the min-entropy until the first non-analyticity is effectively linear, we expect this to capture the leading order of the behavior. As motivated by the free case [27], we expect later non-analyticities to smear out in the thermodynamic limit. For a fixed interaction strength $V=0.8$ (as shown in Fig. 5), we select the $x$-coordinates of the first non-analytic points for different system sizes that we denote as $1-m(t_{c})/M$ and perform a linear regression as shown in Fig. 10. The system parameters treated and the results are summarized in Table 3 and a CSV file containing a summary of regression analyses is provided in the research data management [38]. We find an acute sensitivity on the coupling strength $g$ between the system and the environment, where we find that the non-vanishing critical value persists for higher interaction strengths when the coupling is weaker. For both cases of the coupling ($g=0.5$ and $g=0.25$), the value corresponding to the free case $V=0.0$ indeed matches the value obtained in the exactly solvable free model as carried out in Ref. [27]. At finite $V>0$, for the smaller value of $g$ we find a non-vanishing $t_{c}$-value in the TL for $V\lesssim 1$. However, when further increasing $V$, or in the case of $g=0.5$, we always obtain $t_{c}\approx 0$; a conservative estimate of the numerical errors after the extrapolation gives an errorbar of $\sim 3$ times the standard deviation shown, so that we conclude that these values are zero within our obtained numerical accuracy. This is also true at $V=0$, in agreement with Ref. [27] (see Table 3). Note that we have performed regression analyses for the critical value of decay fraction instead of critical time (as plotted in Fig. 3 for the vNEE); the scaling analyses for the critical time is provided in the Appendix B where we observe the same pattern.
@@
We can make this interpretation concrete to show the vanishing of the critical values in the thermodynamic limit. Consider the strongly interacting case where $V/t_{h}\gg 1$ (where $t_{h}\equiv t_{s}=t_{e}$ which has been set to $1.0$ in this work). Contrary to intuition, repulsive interactions don’t generically lead to particles spreading out. The concept of pairing induced by repulsive interactions is known in the literature [39, 40, 41, 42, 43, 44] where repulsively bound pairs form as a consequence of repulsive interactions . In fact this behavior is prevalent in our setup too where a strong value of $V$, say $V=5.0$ given other parameters are same as in our paper ($g=0.5$, $t_{s}=t_{e}=1.0$, $L=50$ and $M$ can be anywhere between $3$ and $7$ sites) causes the system to freeze and no particles leak into the environment. See Appendix C for $V=3.0$ and $V=5.0$ for time scales accessible to us. Accordingly, in order to write down an ansatz for the number of particles in the system, we have two possibilities: (1) one particle decays to the environment and accordingly a hole enters the system, causing dynamics in the setup  [^1] In this strongly interacting case, no higher order processes will play any relevant role and the system will freeze in this state and no further decay of particles will be observed., and (2) no particle leaks into the environment, thereby having no dynamics at all (see Appendix C). Since we are interested in finding the crossing of the ground state of the entanglement Hamiltonian and the dynamics is dominated by these two possibilities, only the largest and the second-largest Schmidt coefficients, namely $1-\lambda(t)$ and $\lambda(t)$ respectively, are playing a role such that $\lambda(t=0)=0$ (see Appendix C where we have also provided numerical evidences for Schmidt values). The ansatz reads as
@@ VII Summary and Outlook
Along the same lines as in the free case [27], we expect the later non-analyticites (after the first one) to smear out in the TL. Using the linear growth of the min-entropy up to the first non-analyticity, we used linear regression to extrapolate to the thermodynamic limit where we found that there exists a non-zero critical time for weak interactions at which the quantum phase transition in the entanglement Hamiltonian develops. However, as the interaction strength increases, the critical value vanishes in the thermodynamic limit, implying an instantaneous dynamics leading up to the non-analytic behavior (see Fig 12). Note that also for $t_{c}=0$ the semiclassical description loses its applicability, so that, as long as $t_{c}\lesssim{t_{\text{Page }}}$, the proposed scenario keeps its validity (see Fig. 12). Our analyses show that one must be careful before naively applying perturbation theory by switching on the interactions because of the underlying quantum critical point that might influence the behavior of the system even for finite Rényi entropies. Moreover, the order parameter for the transition, whose schematic is sketched in Fig. 7, is the change in particle number density where a depletion transition drives the quantum phase transition. Near the critical point, the particle density develops a power law behavior with an exponent $1/2$ for the free case [27]. Since the weak interactions are nearly the same as the free case, including the thermodynamic critical value (see Table 3), we strongly expect the power law behavior to stay the same. Indeed, as shown in details in Section V, we do find the same value of the critical exponent thereby concluding that such temporal quantum phase transitions are robust despite the presence of interactions which makes our setup non-integrable.
Our work is based on a weak coupling $g=0.50$ between the system in the environment but we found an interesting behavior for further weaker coupling $g=0.25$. Namely, we found an acute sensitivity on the coupling strength $g$ between the system and the environment where the weaker the coupling $g$, the stronger the interaction is required for the thermodynamic critical value to vanish. Furthermore, the critical exponent’s analysis and the finite-size scaling performed in Section V remain no longer valid for $g=0.25$. We do not find a universal collapse and therefore, no associated exponents. This might suggest that the nature of phase transition has changed from a continuous phase transition to a first-order transition at some critical value $g^{\star}$, or something completely different altogether. A true understanding for significantly smaller values of $g$ is lacking, in the sense that why it shows a finite, non-zero value of $t_{c}$ in the thermodynamic limit and why the exponents are no longer universal, whether it be a first-order phase transition or something completely different. This is left as an outlook.
@@
Returning to our original motivation from the black hole information paradox, we have drawn analogies between the black hole setup and our lattice setup. Despite the obvious differences, there does exist some similarities between the two. A recent proposal in black hole physics has gathered attention, namely the island formula [47, 48, 49], where a quantum extremal surface develops during the temporal evolution of vNEE right before the Page time. The presence of islands is crucial after the Page time for the reproducibility of the Page curve. However, the emergence of islands is not smooth, thereby leading to a non-analytic behavior in the evolution of vNEE. This suggests the possibility of whether Page curves in the black hole information paradox can be understood via the mechanism highlighted in this work. Ours is a $1+1$ dimensional setup while black holes are $3+1$ dimensional objects. In one spatial dimension, we find non-analyticity in the ground state energy of the entanglement Hamiltonian at zero fictitious temperature associated with the entanglement Hamiltonian. The possibility we are talking about is that when one goes to higher dimensions, the criticality might not just be limited to zero fictitious temperature but expand to higher, non-zero fictitious temperatures, causing non-analyticity in not just the ground state but excited states of the entanglement Hamiltonian. This includes the von Neumann entropy that might have a non-analyticity in higher dimensions when approached via our methodology as sketched in this work. Numerics become practically impossible in higher dimensions, so the hope rests on analytical frameworks revolving around the entanglement Hamiltonian associated with a Schwarzchild black hole in AdS spacetime (prepared in a pure state as is normally the setup found in the literature, see Ref. [24]) coupled to an empty environment. The question remains about the validity of our claims to black holes, and to which extent we can further understand the emergence of these islands as being related to some form of quantum phase transition in the associated entanglement (equivalently, modular) Hamiltonian.